Mary and her younger brother Alex decide to ride the carousel at the State Fair. Mary sits on one of the horses in the outer section at a distance of 2.0 m from the center. Alex decides to play it safe and chooses to sit in the inner section at a distance of 1.4 m from the center. The carousel takes 5.0 s to make each complete revolution.

(a) What is Mary's angular speed ωM and tangential speed vM?

ωM = rev/s
vM = m/s


(b) What is Alex's angular speed ωA and tangential speed vA?

ωA = rev/s
vA = m/s
Answer #1

a)  

= 0.2 rev/s

= 0.2 × 2π rad/s = 1.256 rad/s

= 1.256 × 2 = 2.51 m /s

b)

= 0.2 rev/s

= 0.2 × 2π rad/s = 1.256 rad/s

= 1.256 × 1.4 = 1.76 m/s